# A DEIM-CUR factorization with iterative SVDs11footnotemark: 1

### v2

@@ 1 Introduction
In data analysis applications and machine learning, the data set is often represented by a matrix $A\in\mathbb{R}^{m\times n}$, which is usually large. In many cases, a key step in the analysis is to approximate the data using a few features and/or a few data points so that one can easily manipulate, understand, and interpret the data. The optimal approximation is obtained by the truncated singular value decomposition (TSVD). On the other hand, this approximation problem can also be reduced to identifying a good subset of columns and rows in the data matrix, a CUR factorization. A CUR decomposition is an alternative solution to the SVD, motivated by the fact that in several applications, it can be challenging to have a concrete interpretation of the singular vectors. Additionally, the singular vectors fail to preserve properties such as nonnegativity, and sparsity, if the original data matrix has these. Theoretical computer science and numerical linear algebra communities have extensively studied column subset selection (index selection) algorithms. In the latter, researchers have primarily focused on deterministic algorithms, which exploit the SVD or rank-revealing QR factorizations that select columns by pivoting rules [3, 7, 22, 33, 21].
On the other hand, researchers in the theoretical computer science community have predominantly directed their attention towards selecting optimal columns using randomized algorithms with provable error bounds [5, 12, 16, 18, 23]. Randomized algorithms sometimes offer reduced computational complexity compared to deterministic methods. Nevertheless, these techniques usually necessitate oversampling columns and rows beyond the desired target rank $k$ to achieve strong provable approximation guarantees.
@@
where the full-rank matrices $C=AP$ and $R=S^{\top}A$ consist of a subset of the columns and rows of A, respectively, and the middle matrix $M$ of full rank is computed such that the decomposition is as close to $A$ as possible. For a general overview on computing the middle matrix, see [25, 26]. Given $C$ and $R$, a standard procedure to determine $M$ is (see, e.g., [29, Sec. 2], where also an alternative is presented) by two consecutive least squares problems:
@@
In this paper, we explore iterative subselection variants of the DEIM scheme. Our contribution aims to improve the approximation quality of the DEIM scheme by iteratively invoking it, in the sense that we select subsequent columns and rows based on the previously selected ones. Thus, we modify $A$ after each iteration by removing the information that has been captured by the previously selected columns and rows. We show this by adapting an existing adaptive sampling technique for the DEIM scheme and also propose a new strategy. We also discuss how iterative procedures for computing a few singular vectors of large data matrices can be used with our proposed methods. To the best of our knowledge, this is the first deterministic DEIM type algorithm for large-scale data sets.
@@ 2 Volume sampling for column subset selection problem
Data: $A\in\mathbb{R}^{m\times n}$, target rank $k$, # rounds $t$, columns per round $c$
@@ 14      $C=A(\,:,\mathbf{p})$;  $E_{i}=A-CC^{+}A$
The volume sampling method of [12] as summarized in Algorithm 2 involves alternating between two steps in each round: selecting a subset of columns and updating the probability distribution over all columns. The selection of columns in each round is influenced by the columns picked in previous rounds. Suppose we aim to select $k$ subsets of columns from matrix $A$. The process begins with an initial probability distribution and randomly selects $c<k$ columns to form a matrix $C$. The selection of columns is based on the norms of the columns, as described in [12, 18]. Each column $j$ is chosen with a probability $\text{pr}_{i}^{(j)}={\lVert E_{i-1}^{(j)}\rVert^{2}}\,/\,{\lVert E_{i-1}\rVert_{F}^{2}}$ (as in Algorithm 2 of Algorithm 2). After selecting $c$ columns, the probabilities are updated based on the chosen columns, and $c$ new columns are sampled and added to the matrix $C$. This iterative process continues until all $k$ columns are selected. Note that assigning zero probability to previously selected indices, as described in Algorithm 2, represents a sampling without replacement strategy.
@@
In this paper, we introduce a modified approach based on the DEIM scheme to implement a variant of the adaptive sampling proposed by Deshpande et al. [12] and propose a new iterative subselection strategy. By incorporating the DEIM scheme, our methods provide a deterministic technique for iterative subselection of column indices, in contrast to the original methods that employ a probabilistic approach [5, 12, 13, 28, 34].
We propose deterministic variants of the adaptive sampling technique for several reasons:
@@
    The volume sampling technique, akin to many randomized algorithms requires oversampling of columns and rows beyond the specified target rank $k$ to achieve strong provable approximation guarantees. In our experience, the deterministic algorithms (like our proposed algorithms) typically yield lower approximation errors compared randomized algorithms when we fix a rank parameter $k$ and choose exactly $k$ columns and rows (see, e.g., Table 4).
@@ 3 Small-scale DEIM type CUR with iterative SVDs
CADP-CUR Fixed number $A-CMR$ Algorithm 5
@@
DADP-CUR Singular value decay-based $A-CMR$ Remark 3.2
@@
The complexity estimates and the other descriptions are similar for the first three algorithms. However, when constructing a CUR factorization using the first two, one needs to perform almost twice the number of SVDs, $X$, and $E$, compared to what is required by Algorithm 5. For the large-scale algorithm, estimating the precise time complexity of computing a low-rank SVD using an iterative method can be challenging since it depends on the total number of inner iterations (denoted by $\text{nr}_{\text{in}}$ in the table) needed.
@@ 3.1 An iterative DEIM with fixed indices per round and one-sided projected residual
We present a deterministic variant of the iterative subselection scheme discussed in Section 2. The newly proposed algorithm (CADP-CX) builds upon the original volume sampling algorithm [12] by leveraging the benefits of the DEIM technique. The procedure is summarized in Algorithm 3. The method involves iteratively selecting a constant number of column indices, denoted as $c$, from $A$ in multiple rounds. We start by computing the leading $c<k$ singular vectors of $A$. Next, we apply the DEIM scheme (Algorithm 1) to these singular vectors, resulting in the first set of $c$ indices. We then update $A$ by computing the residual matrix $E$ using the interpolative decomposition (as described in Algorithm 3 of Algorithm 3). Next, we compute the leading $c$ singular vectors of $E$ and apply the DEIM procedure again to obtain the next set of $c$ indices. This process is repeated until we have selected all $k$ required indices.
@@ 3.2 A new iterative subselection method
In Algorithm 4, we introduce a new iterative subselection strategy (DADP-CX) for a CUR factorization, which differs from the method employed in our proposed Algorithm 3 and the adaptive sampling procedures in previous works [5, 12, 28, 34]. In contrast to the previous strategy, which selects a fixed number of columns or rows in each iteration, this new strategy dynamically adjusts the selection schedule based on the decay of the singular values of the data (the relative magnitudes of the singular values).
@@ 10      $E=A-CX$ (Update matrix)
Subsequently, we proceed to construct an interpolative decomposition using the chosen column indices and compute the residual matrix $E$ by subtracting this approximation from $A$, i.e., $E=A-CC^{+}A$. To determine the next set of indices, we repeat the aforementioned process on $E$. Thus, we compute the leading singular vectors of $E$ corresponding to singular values greater than $\delta$ times the largest singular value of $E$ and repeat the procedure mentioned earlier. We expect that the multiple passes through $A$ would lead to a reduced approximation error. It is worth mentioning that in Algorithm 4, there is no need to compute the initial SVD of $A^{\top}$ since we can store the initial left singular vectors from the SVD of $A$.
In addition to the new selection strategy described in Algorithm 4, we also define an alternative way to compute the residual in the index selection process, which is presented in Algorithm 5. The newly proposed iterative subselection algorithms (Algorithms 3 and 4) and existing adaptive sampling procedures such as those outlined in [5, 12, 13, 28, 34] define the residual as the error incurred by projecting the matrix $A$ onto either the column space of $C$ or the row space of $R$, i.e., $E=A-CC^{+}A$ or $E=A-AR^{+}R$, respectively. In contrast, this new method (DADP-CUR) defines the residual as the error incurred by simultaneously projecting $A$ onto both the column space of $C$ and the row space of $R$. This means computing a CUR factorization at each step using only the selected columns and rows.
@@ 12      $E=A-A(:,\mathbf{p})\cdot M\cdot A(\mathbf{s},:)$ (Update matrix)
Without showing specifics, it is worth mentioning that Algorithm 3 can be adapted using the newly defined residual. Thus, in Algorithm 3, CADP-CUR calculates both the right and left singular values, selects column and row indices in each round, and updates the residual using $E=A-CMR$, where $M=A(:,\mathbf{p})\ \backslash\ (A\ /\ A(\mathbf{s},:))$.
@@ 4 Large-scale DEIM type CUR with iterative SVDs
$E\,(\widehat{V}_{\widehat{k}}Z)=(\widehat{U}_{\widehat{k}}W)\,\widehat{\Sigma},\quad E^{\top}(\widehat{U}_{\widehat{k}}W)=(\widehat{V}_{\widehat{k}}Z)\,\widehat{\Sigma}+\beta_{\widehat{k}}\widehat{\mathbf{v}}_{{\widehat{k}}+1}(W^{\top}\mathbf{e}_{\widehat{k}})^{\top},$
@@
The pair $(\widehat{U}_{k},\widehat{V}_{k})$ may be viewed as a pair of approximate invariant spaces with error $\|\mathbf{f}_{k}\|$. The spaces are expanded with Lanczos bidiagonalization to dimension ${\widehat{k}}$, after which the selection procedure is carried out again. This scheme is repeated until the quantify $\|\mathbf{f}_{k}\|$ is sufficiently small. We summarize the method in Algorithm 6. Note that MATLAB built-in function svds is a different implementation of a related technique.
@@ 7      Partition according to (2), restart with $\widehat{U}_{k}$, $\widehat{V}_{k}$,
Data: $A\in\mathbb{R}^{m\times n}$, desired rank $k$, threshold parameter $\delta\in(0,1]$, max number of indices per round $\ell$
@@ 3       Compute $\sigma_{j}$’s and $\mathbf{v}_{j}$’s by Algorithm 6
4         finding $\sigma_{b}$, the last index $i\leq k-{\sf length}(\mathbf{p})$ with $\sigma_{i}>\delta\,\sigma_{1}$ or
@@ 8       Update an incremental QR decomposition $C=QT$
The efficiency of Algorithm 7 is because for the procedure in Algorithm 6; we do not need the matrix $E$ in explicit form, only matrix-vector products (MVs) with $E$ and $E^{\top}$ are necessary (see Algorithm 7). The routine of Algorithm 6 also takes several MVs that depend on the distribution of the singular value and the starting vector.
The cost of computing the singular values and vectors in Algorithm 7 depends on the total number of inner iterations of Algorithm 6. The number of iterations required by the Krylov–Schur algorithm depends on the size of the matrix. In a matrix-vector product $E\mathbf{x}$ for a vector $\mathbf{x}$, the component $A\mathbf{x}$ costs $\mathcal{O}(mn)$ for a full matrix. In the case of a sparse matrix with $d$ nonzeros entries per row, the cost reduces to $\mathcal{O}(md)$. The aggregated cost of Algorithm 7 is only $\mathcal{O}(mk^{2})$. In Algorithm 7, the computation of $Q(Q^{\top}\mathbf{x})$ requires $\mathcal{O}(mk)$ operations. The cost of solving the least squares problem $M=C^{+}AR^{+}$ would be $\mathcal{O}(mnk)$, which is relatively expensive. Nevertheless, it is important to highlight that this step is necessary for all CUR methods as the final step.
As previously mentioned, it is not necessary to compute the initial SVD of $A^{\top}$ in this case, as we can simply retain the initial left singular vectors obtained from the SVD of $A$. The value of $\delta$ will typically depend on the data set. A value close to $1$ may be favorable for the approximation result but is more expensive since Algorithm 6 needs to be carried out approximately $k$ times. However, having $\delta=1$ implies that we select one index per iteration, and thus, we need just the first right and left singular vectors of $E$ corresponding to the largest singular value. We can reduce the computational cost by specifying an earlier convergence criterion for finding the approximate leading right and left singular vectors. We use Wedin’s theorem for this. The theorem bounds the distance between subspaces and the proof is in (cf., e.g., [30, pp. 260–262]).
@@
Theorem 4.1 shows that the computed singular vectors extracted by the projection method are optimal up to the factor in the right-hand side of the above inequality. This implies that any change in the entries of the computed singular vectors is bounded by this factor. Note that $\Sigma_{2}$ is unknown. For our context, we use $\widehat{\Sigma}_{2}$ as an approximation to $\Sigma_{2}$. Since we are only concerned with approximating the first leading right singular vector $\widehat{\mathbf{v}}_{1}$, we approximate $\alpha\approx\widehat{\sigma}_{1}-\widehat{\sigma}_{2}$. Let $m_{1}(\widehat{\mathbf{v}}_{1})$ and $m_{2}(\widehat{\mathbf{v}}_{1})$ denote the largest and second-largest entries in $\widehat{\mathbf{v}}_{1}$, respectively, and let $\mathbf{f}_{2}=E^{\top}\widehat{\mathbf{u}}_{1}-\widehat{\sigma}_{1}\widehat{\mathbf{v}}_{1}$ be the residual vector (associated with residual matrix $F_{2}$). The above svds routine results in $\mathbf{f}_{1}=E\widehat{\mathbf{v}}_{1}-\widehat{\sigma}_{1}\widehat{\mathbf{u}}_{1}=0$ (associated with residual matrix $F_{1}$). The DEIM algorithm selects the index corresponding to the largest element in the magnitude of a vector. Therefore, when $\delta=1$, one can set an early convergence criterion to find the first singular vector that corresponds to the largest singular value, using the following approximate bound:

### v3

@@ 1 Introduction
In data analysis applications and machine learning, the data set is often represented by a matrix $A\in\mathbb{R}^{m\times n}$, which is usually large. In many cases, a key step in the analysis is to approximate the data using a few features and/or a few data points so that one can easily manipulate, understand, and interpret the data. The optimal approximation is obtained by the truncated singular value decomposition (TSVD). On the other hand, this approximation problem can also be reduced to identifying a good subset of columns and rows in the data matrix, a CUR factorization. A CUR decomposition is an alternative solution to the SVD, motivated by the fact that in several applications, it can be challenging to have a concrete interpretation of the singular vectors. Additionally, the singular vectors fail to preserve properties such as nonnegativity and sparsity, if the original data matrix has these. Theoretical computer science and numerical linear algebra communities have extensively studied column subset selection (index selection) algorithms. In the latter, researchers have primarily focused on deterministic algorithms, which exploit the SVD or rank-revealing QR factorizations that select columns by pivoting rules [3, 7, 21, 31, 20].
On the other hand, researchers in the theoretical computer science community have predominantly directed their attention towards selecting “optimal” columns using randomized algorithms with provable error bounds [5, 12, 16, 18, 22]. Randomized algorithms sometimes offer reduced computational complexity compared to deterministic methods. Nevertheless, these techniques usually necessitate oversampling columns and rows beyond the desired target rank $k$ to achieve strong provable approximation guarantees.
@@
where the full-rank matrices $C=AP$ and $R=S^{\top}A$ consist of a subset of the columns and rows of A, respectively, and the middle matrix $M$ of full rank is computed such that the decomposition is as close to $A$ as possible. For a general overview on computing the middle matrix, see [24, 25]. Given $C$ and $R$, a standard procedure to determine $M$ (see, e.g., [28, Sec. 2], where also an alternative is presented) is by two consecutive least squares problems:
@@
In this paper, we explore iterative subselection variants of the DEIM scheme. Our contribution aims to improve the approximation quality of the DEIM scheme by iteratively invoking it, in the sense that we select subsequent columns and rows based on the previously selected ones. Thus, we modify $A$ after each iteration by removing the information that has been captured by the previously selected columns and rows. We show this by adapting an existing volume sampling technique for the DEIM scheme and also propose a new strategy. We also discuss how iterative procedures for computing a few singular vectors of large data matrices can be used with our proposed methods. To the best of our knowledge, this is the first deterministic DEIM type algorithm for large-scale data sets.
@@ 2 Volume sampling for column subset selection problem
Data: $A\in\mathbb{R}^{m\times n}$, target rank $k$, # rounds of $t$, columns per round $c$
@@ 14      $C=A(\,:,\mathbf{p})$;  $E_{i}=A-CC^{+}A$
The volume sampling method of [12] as summarized in Algorithm 2 involves alternating between two steps in each round: selecting a subset of columns and updating the probability distribution over all columns. The selection of columns in each round is influenced by the columns picked in previous rounds. Suppose we aim to select a subset of $k$ columns from matrix $A$, the process begins with an initial probability distribution and randomly selects $c<k$ columns to form a matrix $C$. The selection of columns is based on the norms of the columns, as described in [12, 18]. Each column $j$ is chosen with a probability $\text{pr}_{i}^{(j)}={\lVert E_{i-1}^{(j)}\rVert^{2}}\,/\,{\lVert E_{i-1}\rVert_{F}^{2}}$ (as in Algorithm 2 of Algorithm 2). After selecting $c$ columns, the probabilities are updated based on the chosen columns, and $c$ new columns are sampled and added to the matrix $C$. This iterative process continues until all $k$ columns are selected. Note that assigning zero probability to previously selected indices, as described in Algorithm 2, represents a sampling without replacement strategy.
@@
In this paper, we introduce a modified approach based on the DEIM scheme to implement a variant of the volume sampling proposed by Deshpande et al. [12] and propose a new iterative subselection strategy. By incorporating the DEIM scheme, our methods provide a deterministic technique for iterative subselection of column indices, in contrast to the original methods that employ a probabilistic approach [5, 12, 13, 27, 32].
We propose deterministic variants of the volume sampling technique for several reasons:
@@
    The volume sampling technique, akin to many randomized algorithms requires oversampling of columns and rows beyond the specified target rank $k$ to achieve strong provable approximation guarantees. In our experience, the deterministic algorithms (including our proposed algorithms) typically yield lower approximation errors compared to randomized algorithms when we fix a rank parameter $k$ and choose exactly $k$ columns and rows (see, e.g., Table 4).
@@ 3 Small-scale DEIM type CUR with iterative SVDs
CADP-CUR Fixed number $A-CMR$ Remark 3.2
@@
DADP-CUR Singular value decay-based $A-CMR$ Algorithm 5
@@
The complexity estimates and the other descriptions in Table 2 are similar for the first four algorithms. However, when constructing a CUR factorization using the first two, one needs to compute almost twice the number of SVDs, $X$, and $E$, compared to what is required by CADP-CUR and DADP-CUR. For the large-scale algorithm, estimating the precise time complexity of computing a low-rank SVD using an iterative method can be challenging since it depends on the total number of inner iterations (denoted by $\text{nr}_{\text{in}}$ in the table) needed.
@@ 3.1 An iterative DEIM with fixed indices per round and one-sided projected residual
We present a deterministic variant of the iterative subselection scheme discussed in Section 2. The proposed algorithm called CADP-CX builds upon the original volume sampling algorithm [12] by leveraging the benefits of the DEIM technique. The method involves iteratively selecting a constant number of column indices, denoted as $c$, from $A$ in multiple rounds. We start by computing the leading $c<k$ singular vectors of $A$. Next, we apply the DEIM scheme (Algorithm 1) to these singular vectors, resulting in the first set of $c$ indices. We then update $A$ by computing the residual matrix $E$ using the interpolative decomposition (as described in Algorithm 3 of Algorithm 3). Next, we compute the leading $c$ singular vectors of $E$ and apply the DEIM procedure again to obtain the next set of $c$ indices. This process is repeated until we have selected all $k$ required indices. The procedure is summarized in Algorithm 3.
@@ 3.2 A new iterative subselection method
In Algorithm 4, we introduce a new iterative subselection strategy referred to as DADP-CX for a CUR factorization, which differs from the method employed in our proposed Algorithm 3 and the adaptive sampling procedures in previous works [5, 12, 27, 32]. In contrast to the previous strategy, which selects a fixed number of columns or rows in each iteration, this new strategy dynamically adjusts the selection schedule based on the decay of the singular values of the data (the relative magnitudes of the singular values).
@@ 10      $E=A-CX$ (Update matrix)
We then proceed to construct an interpolative decomposition using the chosen column indices and compute the residual matrix $E$ by subtracting this approximation from $A$, i.e., $E=A-CC^{+}A$. To determine the next set of indices, we repeat the aforementioned process on $E$. Thus, we compute the leading singular vectors of $E$ corresponding to singular values greater than $\delta$ times the largest singular value of $E$ and repeat the procedure mentioned earlier. We expect that the multiple passes through $A$ would lead to a reduced approximation error. It is worth mentioning that in Algorithm 4, there is no need to compute the initial SVD of $A^{\top}$ since we can store the initial left singular vectors from the SVD of $A$.
In addition to the new selection strategy described in Algorithm 4, we also define an alternative way to compute the residual in the index selection process, which is presented in Algorithm 5. The newly proposed iterative subselection algorithms (Algorithms 3 and 4) and existing adaptive sampling procedures such as those outlined in [5, 12, 13, 27, 32] define the residual as the error incurred by projecting the matrix $A$ onto either the column space of $C$ or the row space of $R$, i.e., $E=A-CC^{+}A$ or $E=A-AR^{+}R$, respectively. In contrast, this new method termed as DADP-CUR, defines the residual as the error incurred by simultaneously projecting $A$ onto both the column space of $C$ and the row space of $R$. This means computing a CUR factorization at each step using only the selected columns and rows.
@@ 12      $E=A-A(:,\mathbf{p})\cdot M\cdot A(\mathbf{s},:)$ (Update matrix)
For the completeness of comparison, in the experiments, we consider a variation of Algorithm 3 referred to as CADP-CUR, where we use the newly defined residual. Thus, in Algorithm 3, CADP-CUR calculates the right and left singular values, selects both column and row indices in each round, and updates the residual using $E=A-CMR$, where $M=C^{+}AR^{+}$.
@@ 4 Large-scale DEIM type CUR with iterative SVDs
$E\,(\widehat{V}_{\widehat{k}}Z)=(\widehat{U}_{\widehat{k}}W)\,\widehat{\Sigma},\quad E^{\top}(\widehat{U}_{\widehat{k}}W)=(\widehat{V}_{\widehat{k}}Z)\,\widehat{\Sigma}+\beta_{\widehat{k}}\widehat{\mathbf{v}}_{{\widehat{k}}+1}(W^{\top}\mathbf{e}_{\widehat{k}})^{\top}.$
@@
The pair $(\widehat{U}_{k},\widehat{V}_{k})$ may be viewed as a pair of approximate invariant spaces with error $\|\mathbf{f}_{k}\|$. The spaces are expanded with Lanczos bidiagonalization to dimension ${\widehat{k}}$, after which the selection procedure is carried out again. This scheme is repeated until the quantity $\|\mathbf{f}_{k}\|$ is sufficiently small. We summarize the method in Algorithm 6. Note that MATLAB built-in function svds is a different implementation of a related technique.
@@ 3       Compute $\sigma_{j}$’s and $\mathbf{v}_{j}$’s by Algorithm 6
4         finding $b$, the last index $i\leq k-{\sf length}(\mathbf{p})$ with $\sigma_{i}>\delta\,\sigma_{1}$ or
@@ 8       Update an incremental QR decomposition $C=QT$
The efficiency of Algorithm 7 is because for the procedure in Algorithm 6 we do not need the matrix $E$ in explicit form; only matrix-vector products (MVs) with $E$ and $E^{\top}$ are necessary (see Algorithm 7). The routine of Algorithm 6 also takes several MVs that depend on the distribution of the singular value and the starting vector. The cost of computing the singular values and vectors in Algorithm 7 depends on the total number of inner iterations of Algorithm 6. The number of iterations required by the Krylov–Schur algorithm depends on the size of the matrix. In a matrix-vector product $E\mathbf{x}$ for a vector $\mathbf{x}$, the component $A\mathbf{x}$ costs $\mathcal{O}(mn)$ for a full matrix. In the case of a sparse matrix with $d$ nonzeros entries per row, the cost reduces to $\mathcal{O}(md)$. The aggregated cost of Algorithm 7 is only $\mathcal{O}(mk^{2})$. In Algorithm 7, the computation of $Q(Q^{\top}\mathbf{y})$ requires $\mathcal{O}(mk)$ operations. The cost of solving the least squares problem $M=C^{+}AR^{+}$ would be $\mathcal{O}(mnk)$, which is relatively expensive. Nevertheless, it is important to highlight that this step is necessary for all CUR methods as the final step.
As previously mentioned, it is not necessary to compute the initial SVD of $A^{\top}$ in this case, as we can simply retain the initial left singular vectors obtained from the SVD of $A$. The value of $\delta$ will typically depend on the data set. A value close to $1$ may be favorable for the approximation result but is more expensive since Algorithm 6 needs to be carried out approximately $k$ times. However, having $\delta=1$ implies that we select one index per iteration, and thus, we need just the first right and left singular vectors of $E$ corresponding to the largest singular value. We can reduce the computational cost by specifying an earlier convergence criterion for finding the approximate leading right and left singular vectors. We use Wedin’s theorem for this. The theorem bounds the distance between subspaces and the proof is in (cf., e.g., [29, pp. 260–262]).
@@
Theorem 4.1 shows that the computed singular vectors extracted by the projection method are optimal up to the factor in the right-hand side of the above inequality. This implies that any change in the entries of the computed singular vectors is bounded by this factor. Note that $\Sigma_{2}$ is unknown. For our context, we use $\widehat{\Sigma}_{2}$ as an approximation to $\Sigma_{2}$. Since we are only concerned with approximating the first leading right singular vector $\widehat{\mathbf{v}}_{1}$, we approximate $\alpha\approx\widehat{\sigma}_{1}-\widehat{\sigma}_{2}$. Let $m_{1}(\widehat{\mathbf{v}}_{1})$ and $m_{2}(\widehat{\mathbf{v}}_{1})$ denote the largest and second-largest entries in $\widehat{\mathbf{v}}_{1}$, respectively, and let $\mathbf{f}_{2}=E^{\top}\widehat{\mathbf{u}}_{1}-\widehat{\sigma}_{1}\widehat{\mathbf{v}}_{1}$ be the residual vector (associated with residual matrix $F_{2}$). The above iterative SVD routine results in $\mathbf{f}_{1}=E\widehat{\mathbf{v}}_{1}-\widehat{\sigma}_{1}\widehat{\mathbf{u}}_{1}=0$ (associated with residual matrix $F_{1}$). The DEIM algorithm selects the index corresponding to the largest element in the magnitude of a vector. Therefore, when $\delta=1$, one can set an early convergence criterion to find the first singular vector that corresponds to the largest singular value, using the following approximate bound: